A patient has an obstruction of her right main bronchus. How would this affect the carbon...

  1. A patient has an obstruction of her right main bronchus. How would this affect the carbon dioxide levels in her right pulmonary veins?

  1. What occupies the dorsal groove between the two lungs?
  1. Given that john has a vital capacity of 4.5L and a residual volume of 1500 mL, what is his total lung capacity? (Please show your work)

  1. During expiration, pressure in the __________ is higher than pressure in the trachea.
  1. Pharynx
  2. Larynx
  3. Esophagus
  4. Bronchi
  1. Given the following, calculate expiratory reserve volume

TV = 450,000 microL

VC = 6.2 L

IRV = 1700 mL
Answer #1

1. Obstruction of right main bronchus causes increase in carbon dioxide levels in right pulmonary veins

2. Dorsal groove between the two lungs contains the Spine (thoracic spine)

3. Total Lung Capacity (TLC) = Vital Capacity (VC) + Residual Volume (RV)

Here it is, TLC = 4.5 L+1500 mL = 6L

4. During expiraion, pressure in the Bronchi is higher than pressure in the trachea

5. Using the Formula, VC = TV+IRV+ERV,

ERV (Expiratory Reserve Volume) = VC - (TV+IRV)

ERV= 6.2L - 2.14L = 4.05L